# Methods for Legal Citation Prediction in the Age of LLMs: An Australian Law Case Study

### v1

@@ 3.1 Results
Pre-training on general text vs. law-specialized text.
The results of the LLM-only experiments (1st panel of Table 1) reveal that pre-training alone is insufficient for achieving satisfactory performance in the citation prediction task. Neither general-domain nor law-specialized models demonstrated reasonable accuracy. For example, Claude Sonnet 3.5 achieved only 15% accuracy in predicting citations when provided with both the query text and the RoC. In contrast, both the 7B and 54B variants of SaulLM performed even worse, achieving 2% accuracy or less, despite being explicitly pre-trained on the Open Australian Legal Corpus (including the NSW Caselaw subset used in this study) and other law-specific datasets.
@@
A significant performance boost is observed when comparing pre-trained LLMs to their instruction-tuned counterparts on our training data. For example, the SaulLM-7B model’s performance jumps from 0% to 51.7% in the more
Figure 2: Average Accuracy and Number of Unique Cases for various citation frequency buckets. Results are for Cite-LLaMA-3.1-8B.
challenging Text-only query setup. Similarly, LLaMA-3.1-8B, a general-purpose model, achieves 46.2% accuracy after instruction fine-tuning, surpassing its much larger pre-trained 70B counterpart. Notably, between the LLaMA and SaulLM backbones, instruction tuning on the same dataset and for the same number of epochs proves more effective for the domain-specific SaulLM. This underscores the critical importance of not only domain-specialized pre-training but also targeted fine-tuning on task-specific data and requirements. As anticipated, Figure 2 demonstrates a negative correlation between predictive accuracy and the citation frequency in the dataset. Cases cited more than 100 times achieve 100% accuracy, while accuracy drops below 40% for cases cited 20 times or fewer. This reflects the challenge of accurately predicting less frequently cited cases.
@@
When comparing the best results of each block that involves retrieval: red rows (corresponding to the generic text-embedding from OpenAI) vs. orange rows (corresponding to embeddings trained exclusively on the Open Australian Legal Corpus) in each block of results in Table 1, it is important to to acknowledge that this comparison is not entirely fair due to differences in embedding dimensionality (3072 vs. 384), the volume of data used, and the training algorithms employed. Nonetheless, we observe in each block exhibiting an intuitive pattern in favour of the domain-specialised AusLaw-embeddings (while being 8×8$\times$8 × smaller in dimensionality). This highlights the effectiveness of domain-specialized embeddings and points to a promising future direction: training larger embeddings tailored to the Australian legal domain.
@@
The results indicate the significant role the granularity of DB plays in the accuracy of the retrieval system. Contrary to our expectations, Catchwords proved to be the least effective granularity, performing worse than Full Cases. By a substantial margin, the best performance was achieved with the RoC Aggregations granularity. This pattern is consistent for both the generic and law-specialised embeddings, and across all setups except for the Retrieval-only were Full Case leads to slightly better results compared with Catchwords.
@@
Comparing all methods involving a retrieval component, Hybrid methods are consistently better than Retrieval-only. Among the Hybrid methods, the Voting Ensemble is the best, followed by RAG and then Query Expansion. The superiority of Voting Ensemble highlights the advantages of combining the predictive quality of instruction-tuned LLMs with the robustness of a retrieval system. In contrast, RAG relies on query augmentation to guide the LLM’s predictions within the context of the Top-5 retrieved citations, while the Query Expansion method instead focuses on re-adjusting the query’s semantic space before searching the retrieval space.
Best configurations, and what we can learn from ACC@5.
The best result across all settings is achieved via Cite-SaulLM-7B model (i.e., the first row highlighted in green). The second best results are under the Cite-LLaMA-3.1-8B model (both in the LLM-only and the Voting Ensemble setups). While the Query expansion results are not competitive at ACC@1, the promising 60.4 for Cite-SaulLM-7B for ACC@5 (the second row highlighted in red) suggests that with improved re-ranking of the Top-5 hits, there is potential to boost accuracy by an additional 10%. We leave further exploration of this to future work.
@@ 4 Conclusion
In this paper, we focused on an Australian law case study for the citation prediction task. We examined various approaches to leveraging Large Language Models and retrieval systems, evaluating their effectiveness both independently and in combination. Our findings demonstrate that while pre-training large language models on general or even domain-specialized legal texts is a necessary starting point, it is far from sufficient for achieving satisfactory citation prediction accuracy in the Australian legal domain. The most contributing factor for improving performance lies in targeted instruction tuning with task-specific data, which dramatically boosts accuracy. Our experiments further reveal the importance of choosing the right embedding model and database granularity for retrieval, with results showing up to 70% variation in performance under different granularities. Among the retrieval-augmented methods, ensemble voting strategy stands out as the most effective, outperforming methods like RAG and query expansion.
As potential for future explorations, training better embeddings (i.e., larger dimensionalities, with specific attention to discriminative features), and re-ranking methods to better utilise the Top-k outputs from retrieval system are promising avenues. Additionally, in this paper we did not test the methods in the out-of-distribution setting which could further highlight other potential challenges.
@@ A.1 Three Examples of Catchwords
A.2 Details of Instruction Fine-tuning
The instruction details and training configurations used for instruction fine-tuning are listed in Table 2.
@@ A.2 Details of Instruction Fine-tuning
A.3 Prompt details for various experiments
All prompting details (or few-shot demonstrations when applicable) used in our experiments are provided in Table A.3-A.3.
@@ A.3 Prompt details for various experiments
Prompt template and the few-shot demonstrations used in the LLM+Retrieval experiments of Table 1
@@
Table 6: The prompt template used for re-ranking with LLMs in the RAG experiment of - corresponding to Catch Words - in Table 1. Similar prompts are used for RoC Aggregations, and Full Cases experiments.

### v2

@@ 5.1 Results
Pre-training on general text vs. law-specialised text.
The results of the LLM-only experiments (1st panel of Table 1) reveal that pre-training alone is insufficient for achieving satisfactory performance in the citation prediction task. Neither general-domain nor law-specialised models demonstrated reasonable accuracy. For example, Claude Sonnet 3.5 achieved only 15% accuracy in predicting citations when provided with both the query text and the RoC. In contrast, both the 7B and 54B variants of SaulLM performed even worse, achieving 2% accuracy or less, despite being explicitly pre-trained on the Open Australian Legal Corpus (including the NSW Caselaw subset used in this study) and other law-specific datasets. This highlights that mere domain-specific pre-training of relatively large (i.e., SaulLM-54B) is not sufficient for fine-grained tasks.
Also it is noteworthy that out of the 94-520B tokens used in pre-training of such Lawps-specialised LLMs, only 0.5B tokens were covering Australian jurisdiction (see Table 1s of Colombo et al. (2024b) and Colombo et al. (2024a)). This opens a natural question: whether specific pre-training solely on that segment of 0.5B tokens will give a more jurisdictional chance to the LLM in our task. While a proper investigation of this will require a substantial GPU support, we investigate this question for the 7B-8B LLM scales in Subsection 5.2.
@@
A significant performance boost is observed when comparing pre-trained LLMs to their instruction-tuned counterparts on our training data. For example, the SaulLM-7B model’s performance jumps from 0% to 51.7% in the more challenging Text-only query setup. Similarly, LLaMA-3.1-8B, a general-purpose model, achieves 46.2% accuracy after instruction fine-tuning, surpassing its much larger pre-trained 70B counterpart. Notably, between the LLaMA and SaulLM backbones, instruction tuning on the same dataset and for the same number of epochs proves more effective for the domain-specific SaulLM. This underscores the critical importance of not only domain-specialised pre-training but also targeted fine-tuning on task-specific data and requirements.
As anticipated, Figure 1 demonstrates a negative correlation between predictive accuracy and the citation frequency in the dataset. Cases cited more than 100 times achieve 100% accuracy, while accuracy drops below 40% for cases cited 20 times or fewer. This reflects the challenge fine-tuning faces in accurately predicting less frequently cited cases (i.e., freq<20), where alternative approaches which integrate retrieval (i.e., retrieval with re-ranking, explored next) might offer a more reliable solution.
Figure 1: Average Accuracy and Number of Unique Cases for various citation frequency buckets. Accuracies per bin are based on the following settings from Table 1: Cite-SaulLM-7B (ACC@1: 51.7), RAG (ACC@1: 42.9), and Re-ranker (ACC@1: 52.1), and from Table 2: Cite-AusLawLLM-7B (ACC@1: 52.0).
@@
When comparing the best results of each block that involves retrieval: red rows (corresponding to the generic text-embedding from OpenAI) vs. orange rows (corresponding to embeddings trained exclusively on the Open Australian Legal Corpus) in each block of results in Table 1, it is important to to acknowledge that this comparison is not entirely fair due to differences in embedding dimensionality (3072 vs. 384), the volume of data used, and the training algorithms employed. Nonetheless, we observe an overall intuitive pattern in favour of the domain-specialised AusLaw-embeddings, or at least being on-par with OpenAIs emeddings (while being 8×8$\times$8 × smaller in dimensionality). This highlights the benefits of domain-specialised embeddings and suggests a few promising future directions to explore: training larger embeddings tailored to the Australian legal domain, while we do not investigate this further, a trivial direction of exploration could be to build higher dimensional representations, as well as improving the geometric utilisations of the representation space (e.g., through isotropy Liu et al. (2021)). We leave further explorations of these to future work.
@@
Probing the results along the Index Granularity dimension indicates the significant role the granularity of database (index) plays in the accuracy of the retrieval system. Contrary to our expectations, Catchwords proved to be the least effective granularity, performing worse than Full Cases. By a substantial margin, the best performance was achieved with the RoC Aggregations granularity. This pattern is consistent for both the generic and law-specialised embeddings, and across all setups except for the Retrieval-only were Full Case leads to slightly better results compared with Catchwords.
@@
Comparing all retrieval-based methods, excluding re-ranking, hybrid methods are consistently better than Retrieval-only. The Voting Ensemble is the best, followed by RAG and then Query Expansion. The superiority of Voting Ensemble highlights the advantages of combining the predictive quality of instruction-tuned LLMs with the robustness of a retrieval system. In contrast, RAG relies on query augmentation to guide the LLM’s predictions within the context of the Top-5 retrieved citations, while the Query Expansion method instead focuses on re-adjusting the query’s semantic space before searching the retrieval space.
What we can learn from ACC@5.
The general pattern across all experiments suggests the potential that lies within Top-5 retrieved citations. For instance, while the Query expansion results are not competitive at ACC@1, the promising 60.4% for Cite-SaulLM-7B for ACC@5 (the second row highlighted in red) suggests that with improved re-ranking of the Top-5 hits, there is potential to boost accuracy by up to 10%. We explore this next.
Re-ranker: the Gap of ACC@1 and ACC@5.
The ACC@5 results indicate that the correct citation is often present within the top five retrieved candidates. Building on this insight, the re-ranker’s goal is to enhance the likelihood of selecting the correct citation from these Top-5 candidates. The bottom two blocks in Table 1 demonstrate the performance of various re-ranking methods. Notably, the best-performing re-ranker boosts ACC@1 from 35.1% to 52.1%, significantly narrowing the gap toward the upper bound defined by ACC@5 (60.4%).
5.2 Australian Law Pre-training
SaulLM-7B was pre-trained on a 94B-token general-domain legal corpus comprising various legal sources, including 0.5B tokens from the Open Australian Legal Corpus. To investigate whether pre-training solely on Australian legal data is sufficient for building an effective Australia-specific legal LLM, we pre-train the same underlying vanilla Mistral-7B model and a LLaMA-3.1-8B model (used in SaulLM), on the 0.5B tokens of Australian Legal corpus. Following pre-training, we apply the same citation instruction-tuning procedure on the citation prediction task. While we only pre-train the underlying LLM for 5 epochs on the 0.5B tokens, due to infrastructure limitations, we observe promising benefits of jurisdictional pre-training. The results, presented in Table 2, show that Cite-AusLawLLM-7B slightly outperforms Cite-SaulLM-7B (both built on Mistral-7B), despite the latter being trained on a much larger 94B-token corpus. Furthermore, we observe that pre-training significantly boosts performance of the LLaMA-3.1-8B model after Australian Law pre-training pre-training (Cite-AusLawLLM-8B).
Model Law Data Accuracy
Cite-SaulLM-7B (Mixed Law Pre-training) 94B 51.7
Cite-AusLawLLM-7B (AusLaw Pre-training) 0.5B 52.0
Cite-LLaMA-3.1-8B (No Law Pre-training) - 46.2
Cite-AusLawLLM-8B (AusLaw Pre-training) 0.5B 50.0
Table 2: The results of different pre-training data (Mixed vs. AusLaw-specific) before doing instruction-tuning on citation prediction task. Data size is reported in tokens.
@@ 6 Conclusion
In this paper, we proposed a new benchmark, AusLaw Citation Benchmark, for legal citation prediction and examined various approaches of leveraging Large Language Models and retrieval systems, evaluating their effectiveness both independently and in combination. Our findings demonstrate that while pre-training large language models on general or even domain-specialised legal texts is a necessary starting point, it is far from sufficient for achieving satisfactory citation prediction accuracy in the Australian legal domain. The most contributing factor for improving performance lies in targeted instruction tuning with task-specific data, which dramatically boosts accuracy. Our experiments further reveal the importance of choosing the right embedding model and database granularity for retrieval, with results showing up to 70% variation in performance under different granularities. Among the retrieval-augmented methods, ensemble voting strategy stands out as the most effective, outperforming methods like RAG and query expansion. Furthermore, we show that training re-rankers effectively harnesses the untapped gains available in top-k retrieval accuracy.
While the above approaches offer varying degree of effectiveness, there remains a noticeable $50\%$ gap in performance that calls for further developments. We hope the proposed benchmark and the investigated approaches provide a reliable framework for development of more advanced solutions that intersect LLMs, fine-tuning, pre-training, and retrieval mechanisms.
Limitations
First, due to computational constraints, we limit most of our open-source experiments to 7B-parameter models and do not evaluate the performance of larger-scale models (i.e., SaulLM 141B), which may offer additional gains. Second, our citation prediction task is currently framed as a single-citation prediction problem, where each input maps to only one citation. A more realistic and challenging setting would involve predicting multiple citations per instance under a larger context. Additionally, in this paper we did not test the methods in the out-of-distribution setting which could further highlight other potential challenges in legal domain.
@@ A.1 Three Examples of Catchwords
A.2 Distribution of Citations
Figure 3 (Top) shows the frequency distribution of all $18,677$ citations in the data, and Figure 3 (Bottom) shows the top-20 most frequent citations.
A.3 Details of Instruction Fine-tuning
The instruction details and training configurations used for instruction fine-tuning are listed in Table 3 and Table 10. The re-rankers use the same training configurations.
@@ A.3 Details of Instruction Fine-tuning
A.4 Prompt details for various experiments
All prompting details (or few-shot demonstrations when applicable) used in our experiments are provided in Table A.4-A.4.
@@ A.4 Prompt details for various experiments
[…]
@@
Table 7: The prompt template used for re-ranking with LLMs in the RAG experiment—corresponding to Catch Words—in Table 1. Similar prompts are used for RoC Aggregations and Full Cases experiments.
Prompt template for merging the citation reasons of each case in re-ranker setting 2
system_prompt = """ You will be provided with a case from NSW Case Law, along with reasons it may be cited. Your task is to concisely summarize the reasons into a few sentences (2~3 sentences), if the original citation reasons are too long. """ prompt = """ Casename: CASENAME Citation Reasons: CITATIONREASON """
Table 8: The prompt template for merging the citation reasons of each case into text of similar length.
The prompt template used for generating the citation rationale in re-ranker setting 3
ranker_reason_sys_prompt = """ The following description belongs to a case in the NSW Case Law, you should predict the citation in the text. You will be given a brief text, 5 potential citations, their corresponding citation reasons and a ground-truth correct citation with its potential citation reasons. Note that the ground-truth citation may not be included among the five potential options. Your task is to predict the missing citation, provide a rationale before you draw a conclusion. Your predicted citation should be the ground-truth correct citation, but dont let on that you already know the answer (Dont mention anything about the ground-truth). The rationale should not be over 2~3 sentences. """ ranker_reason_prompt = """ Text: TEXT Potential Citations: CITATION1 Citation Reasons: CITATIONREASON1 CITATION2 Citation Reasons: CITATIONREASON2 CITATION3 Citation Reasons: CITATIONREASON3 CITATION4 Citation Reasons: CITATIONREASON4 CITATION5 Citation Reasons: CITATIONREASON5 Ground-truth Correct Citation: GROUNDTRUTH Ground-truth Citation Reasons: GTREASON """
Table 9: The prompt template used for generating the ground-truth citation rationale given candidate citations in re-ranker setting 3.
Instruction Predict the citation in the text.
Setting 1 Input
test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’) +
Potential Citations:
CITATION1
Citation Reasons: FIRST_CITATIONREASON1
CITATION2
Citation Reasons: FIRST_CITATIONREASON2
CITATION3
Citation Reasons: FIRST_CITATIONREASON3
CITATION4
Citation Reasons: FIRST_CITATIONREASON4
CITATION5
Citation Reasons: FIRST_CITATIONREASON5
Output <GOLD CITATION>
Setting 2 Input
[…]
Potential Citations:
CITATION1
Citation Reasons: MERGED_CITATIONREASON1
CITATION2
Citation Reasons: MERGED_CITATIONREASON2
CITATION3
Citation Reasons: MERGED_CITATIONREASON3
CITATION4
Citation Reasons: MERGED_CITATIONREASON4
CITATION5
Citation Reasons: MERGED_CITATIONREASON5
Output GOLD_CITATION_REASON + <GOLD CITATION>
Setting 3 Input
test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’) +
Potential Citations:
CITATION1
Citation Reasons: MERGED_CITATIONREASON1
CITATION2
Citation Reasons: MERGED_CITATIONREASON2
CITATION3
Citation Reasons: MERGED_CITATIONREASON3
CITATION4
Citation Reasons: MERGED_CITATIONREASON4
CITATION5
Citation Reasons: MERGED_CITATIONREASON5
Output RATIONAL + <GOLD CITATION>
Setting 4v1 Input
test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’) +
Potential Citations:
CITATION1
Citation Reasons: Top-1_CITATIONREASON1
CITATION2
Citation Reasons: Top-1_CITATIONREASON2
CITATION3
Citation Reasons: Top-1_CITATIONREASON3
CITATION4
Citation Reasons: Top-1_CITATIONREASON4
CITATION5
Citation Reasons: Top-1_CITATIONREASON5
Output RETRIEVED_TOP_1_CITATION_REASON + <GOLD CITATION>
Setting 4v2 Input
test_set[i][’citation_text’].replace(test_set[i][’cited_case_name’], ’<CASENAME>’) +
Potential Citations:
CITATION1
Citation Reasons: Top-1_CITATIONREASON1
CITATION2
Citation Reasons: Top-1_CITATIONREASON2
CITATION3
Citation Reasons: Top-1_CITATIONREASON3
CITATION4
Citation Reasons: Top-1_CITATIONREASON4
CITATION5
Citation Reasons: Top-1_CITATIONREASON5
Output GOLD_CITATION_REASON + <GOLD CITATION>
Table 10: The used instruction, input and output for the training of different re-rankers.